Mark is 64 years younger than Chasity. 4 years ago, Chasity's age was 3 times Mark's age. How old is Mark now?

Answer by ikleyn(53765)

.
Let M be the Mark's age.


Then the Chasity age is (M+64) years, according to the condition.


4 years ago:


    (M+64) - 4 = 3(M-4)


    M + 60 = 3M - 12


    72 = 3M - M

    72 = 2M  ====>  M = 72/2 = 36.


Answer.  Mark is 36 years now.

Solved.